OpenStudy (anonymous):

What is the discontinuity and zero of the function f(x) = 3x^2+x-4/x-1 ? a. Discontinuity at (-1, 1), zero at ( 4/3 , 0) b. Discontinuity at (-1, 1), zero at ( -4/3 , 0) c. Discontinuity at (1, 7), zero at ( 4/3 , 0) d. Discontinuity at (1, 7), zero at ( -4/3 , 0)

OpenStudy (anonymous):

Set the denominator equal to zero and solve in order to find the x location of the discontinuity. Plug that value back in to the function in order to find the y location.

OpenStudy (shamil98):

f(x) = 3x^2+x-4/x-1 = (3x+4)(x-1)/(x-1) = 3x + 4 3x + 4 =0 3x = -4 x = -4/3 that is your zero.

OpenStudy (shamil98):

The discontinuity. Like euler has said is simply x - 1 =0 x = 1 so. (1,-1), and (-4/3,0)

OpenStudy (anonymous):

The function to plug the discontinuity into is the one that shamil98 found after factoring and simplifying... 3x+4. This returns 7 for the y value. If you had plugged that into the original, you would get 0/0 (whoops, minor oversight ;) )